OpenStudy (anonymous):

What are two numbers that when multiplied is -84 and when added is 5?

OpenStudy (inkyvoyd):

ab=-84 a+b=5

OpenStudy (inkyvoyd):

why don't we use substitution? rewrite the the second equation. ab=-84 b=5-a

OpenStudy (unklerhaukus):

a=5-b (5-b)b=-84

OpenStudy (inkyvoyd):

trololol

OpenStudy (anonymous):

trolling is not appreciated. I'm actually really confused >_<

OpenStudy (unklerhaukus):

\[(5-b)b=-84\] \[5b-b^2+84=0\]

OpenStudy (anonymous):

WHAT